OpenStudy (andriod09):

Comtinuing the previous question: My lawyer and I realized that 6 of those 200 minutes that he had billed me for last month were minutes spent talking about fishing and not law stuff. Here is this month's bill. Fill in the blanks. (bill in comments)

OpenStudy (andriod09):

Bill: Law bill last month's bill 600 Payemnt recieved 600 Amount overdue 0 This month's work 430 minutes x ($3) 1290 Correction for last month ? x($3) -18 Total amount owing ?

OpenStudy (andriod09):

@.UserNotFound. @jazy, @skler, @Bailey_Reona @headge @evergirl @ganeshie8 @annas @Kaitlyn_The_Math_wiz

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

sorry i suck at math

OpenStudy (andriod09):

okay.

OpenStudy (andriod09):

i need help

OpenStudy (anonymous):

sory not good a math

OpenStudy (andriod09):

@seashell

OpenStudy (andriod09):

can some one please help me?

OpenStudy (anonymous):

wait... where are the blanks?

OpenStudy (andriod09):

the ?s i need to find them out

hartnn (hartnn):

ok, since ? * 3 =-18 u can find first ? as \(\huge?=-18/3 = ??\)

hartnn (hartnn):

or infact minutes cannot be negative so solve this \(?=18/3=??\)

OpenStudy (andriod09):

minutes aren't negative, thats what the laywer is paying back

hartnn (hartnn):

ok,yes. so what u got first ? as ? 18/3 =

OpenStudy (andriod09):

it would be 18/3=?| where ? was equal to the number of minutes that wern't spent talking about law stuffs.

hartnn (hartnn):

thats what i am telling, isn't it?

OpenStudy (andriod09):

ik i was making sure i had it right.

hartnn (hartnn):

yup, u had it right ^_^

OpenStudy (andriod09):

so whats next?

hartnn (hartnn):

for total amount owing u would add Amount overdue + This month's work + correction for last month. so Amount owing = 0+1290+(-18) = ?

OpenStudy (andriod09):

The equation would be 1290-18=_??

hartnn (hartnn):

yup, thats correct.

OpenStudy (andriod09):

wow. i got this wrong last time and it was sooooo easy. I must fail epicly. 0.o0.o0.o0.o0.o0.o0.o0.o0.o0o.0o.o0.o0.o0.o0.o0o.0o.o0.o0.o0o.0o.0o.o0.o0.o0.o0o.0o.o0.o0.o0o.0o.0o.o0.o0.o0.o0o.o0.o0.oo0.o0.o0.o0.o0.o0.o0o.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o.0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o0.o

OpenStudy (andriod09):

Thanks for your help dude, i fanned you jsuk

hartnn (hartnn):

welcome and thanks :)